OpenStudy (anonymous):

simplify x/2 + x/4 - 2x/3

OpenStudy (blurbendy):

first you need to find the least common denominator between 2, 4, and 3

OpenStudy (mayankdevnani):

\[\huge \frac{x}{2} + \frac{x}{4} - \frac{2x}{3}\] now find the LCM(least common denominator) of 2,4 and 3?

OpenStudy (mayankdevnani):

can you find the LCM ? @Bherly

OpenStudy (anonymous):

make the denominators 12 first

OpenStudy (anonymous):

can u help me guys to find out the solution?

OpenStudy (anonymous):

(6x/12)+(3x/12)-(8x/12)

OpenStudy (anonymous):

@Krishnadas thats the answer?

OpenStudy (anonymous):

no,(6x+3x-8x)/12 =x/12

OpenStudy (mayankdevnani):

 so 12 is the LCM